### v1

# The importance of visual modelling languages in generative software engineering

Abstract
Multimodal GPTs represent a watershed in the interplay between Software Engineering and Generative Artificial Intelligence. GPT-4 accepts image and text inputs, rather than simply natural language. We investigate relevant use cases stemming from these enhanced capabilities of GPT-4. To the best of our knowledge, no other work has investigated similar use cases involving Software Engineering tasks carried out via multimodal GPTs prompted with a mix of diagrams and natural language.
1 Introduction
Software engineering (SE) applies engineering principles to the development, operation, and maintenance of software systems, ensuring they are reliable, efficient, and meet user requirements. Generative Artificial Intelligence (GenAI) models, such as pre-trained models Devlin et al. (2019); Vaswani et al. (2017) and large language models (LLMs), are revolutionising fields like computer vision and natural language processing through their ability to generate novel and contextually-appropriate content. In recent times, the interplay between SE and GenAI has been receiving increasing attention. A vast array of applications of pre-trained models and LLMs are surveyed in Huang et al. (2024); Hou et al. (2024). However, when it comes to the type of data used in existing studies, it appears that all studies surveyed focused on text-based datasets, with the most prevalent type of data utilised in training LLMs for SE tasks bring programming tasks/problems expressed in natural language Hou et al. (2024). Similar conclusions are reached in Huang et al. (2024), which focused on 7 sub-tasks of SE: requirements generation, code generation, code summarisation, test generation, patch generation, code optimisation, and code translation. In all cases, the pipeline typically begins with instructions in natural language that need to be used in the context of one or more of these 7 sub-tasks.
In SE it is often the case that communication among developers, and between developers and customers, occurs in the form of sketches and diagrams Baltes and Diehl (2014) and not just via natural language. The reason for this can be easily understood once we consider the following excerpt taken from Tufte (2003).
A TALK, which proceeds at a pace of 100 to 160 spoken words per minute, is not an especially high-resolution method of data transmission. Rates of transmitting visual evidence can be far higher. The artist Ad Reinhardt said, ”As for a picture, if it isn’t worth a thousand words, the hell with it.” People can quickly look over tables with hundreds of numbers in the financial or sports pages in newspapers. People read 300 to 1,000 printed words a minute, and find their way around a printed map or a 35mm slide displaying 5 to 40 MB in the visual field. Often the visual channel is an intensely high-resolution channel.
It is then not surprising that the ambition of automatically turning sketches and diagrams into working code, or to reverse engineer working code into diagrams, has existed for a very long time in SE. Not only sketches and diagrams represent a high-resolution communication channel, but when they are drawn following a standard, they become a form of technical languages. The importance of technical languages is that they are denotative: they say one thing and one thing only. Conversely, natural language is connotative: the meaning of a statement is context dependent. It is the intrinsic ambiguity of natural language that makes it not well suited for programming, or for communicating programming-related matters.
In this work, we argue that the advent of multimodal GPTs, such as GPT-4 OpenAI (2024), may represent a watershed in the interplay between SE and GenAI. GPT-4 accepts image and text inputs, and it can therefore receive prompts that contain sketches and diagrams, rather than simply natural language. We therefore investigate relevant use cases stemming from these enhanced capabilities of GPT-4. To the best of our knowledge, no other work has investigated similar use cases involving SE tasks carried out via multimodal GPTs prompted with a mix of diagrams and natural language.
The rest of this paper is organised as follows: Section 2 provides relevant background in GenAI and SE. Section 3 surveys contemporary academic literature on role of GenAI in SE. Section 4 investigates the role of multimodal GPTs in software development by illustrating a selection of use cases. Section 5 provides some concluding remarks by reflecting on how the former use cases address several research gaps in Generative Software Engineering (GenSE), as well as some long standing issues in SE.
2 Background
This section provides relevant background in GenAI and SE.
2.1 Software Engineering
The term “software engineering” was coined by Margaret Hamilton, who developed on-board flight software for NASA’s Apollo program. Key milestones in the field include the development of the Waterfall model (1970s), the introduction of Object-Oriented programming (1980s), the Agile Manifesto (2001), which advocates for flexible, iterative development, the rise of DevOps (2000s) and of continuous integration/continuous delivery (CI/CD) pipelines (2010s).
In the original Waterfall model, the software development process proceeds through a series of sequential phases: requirement analysis; software design; software development; software quality assurance; and software maintenance. To address the inflexibility of the Waterfall model, the increasing complexity of software, and poor customer satisfaction, in the early 2000s a set of frameworks and practices were proposed based on the values and principles expressed in the Manifesto for Agile Software Development Beck et al. (2001). Core values and principles of Agile include: favouring individuals and interactions over processes and tools; working software over comprehensive documentation; customer collaboration over contract negotiation; and responding to change over following a plan. Agile methodologies promote iterative development, frequent feedback, and adaptability to changing requirements throughout the aforementioned phases.
At the same time, the Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al. (2004) utilised across the phases of the software development process. UML is a general-purpose standardised visual language for designing systems. In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system. Use case development (or development of user stories in Agile) is a key step of requirement analysis. Use cases describe how a user interacts with a system or product to achieve a specific goal. Static analysis of a software system describes the structure of a system by showing its classes, their attributes, operations (or methods), and the relationships among objects. Dynamic analysis expresses and model the behaviour of the system over time. More specifically, in UML, use case diagrams support use case development. Class diagrams support static analysis; and interaction (sequence, activity, collaboration) diagrams support dynamic analysis. However, it should be noted that UML is a vast language that finds countless applications to go beyond the three areas here considered.
Despite UML being the standard visual language for designing systems, recent studies suggest that that most software developers favour informal hand drawn diagrams and do not use UML; those using UML, tend to use it informally and selectively Baltes and Diehl (2014). The authors further advocated the development of suitable tools to make better use of such sketches. Our study argues that GenAI may represent such missing tool.
2.2 Generative Artificial Intelligence
GenAI refers to a class of AI models that can create new content, such as text, diagrams, or code, based on the patterns they learn from existing data Eapen et al. (2023). Large Language Models (LLMs) are a specific category of GenAI models that focus on language-related tasks, such as text generation, translation, summarisation, and question answering. LLMs are typically based on a neural network architecture called a Transformer Vaswani et al. (2017), which is Pre-trained — hence the name Generative Pretrained Transformer (GPT) — on a massive dataset of text and code, and which allows them to process and generate text by considering long-range dependencies and contextual information Brown et al. (2020). The latest GPTs, such as GPT-4 OpenAI (2024), are multimodal: they accept image and text inputs and produce text outputs, leading to new applications Li et al. (2024b).
2.2.1 Prompt engineering
Prompts are functions that map an example from a dataset to a natural language input and target output Bach et al. (2022). Due to the cost of re-training LLMs, prompt engineering, that is the use of prompts to train and query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning. Key prompt engineering techniques include: Single Prompt Techniques, Multiple Prompt Techniques, and the use of LLM frameworks with external tools.
Single Prompt Techniques include Zero-Shot, Few-Shot, Chain of Thought (CoT), and Program-Aided Language Models (PAL). Zero-Shot Prompting involves providing tasks with natural language without additional context, relying on the model’s pre-existing capabilities. Few-Shot Prompting includes providing examples within the prompt to guide the model, enhancing its performance on complex tasks by exposing it to the input and output patterns Brown et al. (2020). Chain of Thought Prompting Wei et al. (2024) aids in breaking down reasoning tasks into smaller steps to improve processing and outcome accuracy, using either zero-shot or few-shot methods to encourage step-by-step thinking. PAL extends this by integrating code into reasoning processes to produce more accurate outputs Gao et al. (2023).
Multiple Prompt Techniques leverage combining different strategies to refine outputs. Voting/self-consistency Wang et al. (2023) involves generating multiple responses and selecting the most common result, which can improve accuracy, especially for complex reasoning tasks. Divide and Conquer methods split tasks into subtasks handled in sequence for improved manageability and precision, seen in Directional Stimulus Prompting Li et al. (2024a), Generated Knowledge Liu et al. (2021), and Prompt Chaining. Self-evaluation asks the model to verify output accuracy, exemplified by Reflexion Shinn et al. (2024) and Tree of Thoughts Yao et al. (2024), enabling iterative improvement.
Retrieval-Augmented Generation (RAG) Fan et al. (2024) and ReAct Yao et al. (2023) combine LLMs with external systems to improve context handling and output relevance. RAG uses vector search for context retrieval before passing it to an LLM for output generation. ReAct combines reasoning traces and task-specific actions, allowing models to interact with external knowledge sources to refine outputs.
3 The role of GenAI in Software Engineering
In recent times, SE has become one of the important application areas for GenAI. We focus on two recent surveys Huang et al. (2024); Hou et al. (2024) investigating the interplay between SE and GenAI.
Several studies, e.g. Arora et al. (2024); White et al. (2024), investigated the use of GenAI in the context of the requirement engineering sub-task. We do not believe sketches and diagram play a significant role in this sub-task, which is mainly concerned in turning requirements expressed in natural language by the customer into suitable user stories and/or conceptual diagrams Robeer et al. (2016).
SE sub-tasks of interest in the context of the present study include: software design, software development, and code summarisation.
Application of LLMs in software design remains relatively sparse: Huang et al. (2024) does not cover this sub-task, while Hou et al. (2024) only report 4 works Kolthoff et al. (2023); Mandal et al. (2023); White et al. (2024); Zhang et al. (2024), none of which overlaps with the content of the present study; they also stress that by expanding the use of LLMs to this under-explored area, it is possible to improve how software designs are conceptualised.
Conversely, both surveys identify a plethora of works concerned with software development (including code generation, test case generation, patch generation, and code optimisation) and code summarisation.
Code generation has been object of investigation for a long time in the AI community. Early works used symbolic and neural-semiotic approaches Alur et al. (2013). However, recent neurolinguistic models, such as GPT-4 Liu et al. (2024) and Copilot Ma et al. (2023), can generate code directly from natural language descriptions. This capability has significantly advanced code completion, automatic code generation, and the transformation of natural language annotations into code, empowering developers with powerful tools and driving automation in software development. None of the studies listed in the above surveys focus on code generation that leverages multimodal prompts that include UML diagrams.
Code summarisation Ahmed et al. (2024) aims to automatically generate descriptions of a given source code. This technique improves code comprehension, documentation, and collaboration by providing clear summaries. Existing studies in code summarisation focus on analysing code structures and contexts to generate informative natural language summaries. None of the studies listed in the above surveys focus on code summarisation producing a UML diagram as its output.
Finally, while there exist a few studies that investigated the generation of UML diagram with support from LLMs Conrardy and Cabot (2024); Wang et al. (2024); Cámara et al. (2023), none of these studies go as far as investigating the generation of working code from UML diagrams, as well as the reverse engineering of relevant UML diagrams from existing code. Perhaps the most interesting study among the three listed is Cámara et al. (2023), which focus on how to build UML class diagrams by using ChatGPT as a modelling assistant by leveraging the PlantUML language, which we will also adopt in the following sections.
4 Multimodal GPTs in software development
In what follows, we will assume that a preliminary requirement analysis has been carried out, which has produced a portfolio of initial user stories. While GPTs may find applications in the realm of the development of user stories and in the translation of user stories into working software, as things stand today, this level of automation in the realm of code generation does not exist in SE practice. Conversely, what we find in practice are teams of software developers tasked with translating user stories into working software. Our focus in this section, is to explore novel use cases of GPTs in this specific context.
The role of sketches and diagrams in the daily work of software developers has been investigated in Baltes and Diehl (2014). This study found that most practitioners produced informal hand drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively. We argue that this stems from the fact that the adoption of a “formal” notation, at present, bears no advantage with respect to an “informal” one. Over several decades, firms have repeatedly tried to develop tools (e.g. IBM Rational Rose) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code. While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools. We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one for generating complete code. In other words, UML diagrams are most useful when they are high level and sufficiently abstract. Generation of complete working code requires such diagrams to reach a level of detail that is equivalent to the working code itself; but generating diagrams at this level of abstraction would be a complete waste of time: why then not generating the code itself directly?
Albeit we believe there is a good reason for practitioners to produce informal hand drawn diagrams, we also believe that this practice is set to change with the advent of multimodal GPTs, such as GPT-4 OpenAI (2024). In what follows, we will outline a set of use cases illustrating how practitioners may combine multimodal GPTs and UML diagrams to innovate SE practices. We will utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
4.1 Static modelling
As a motivating example to illustrate our use cases, we consider the introductory case study in Section 3 of Stevens and Pooley (2006). In this section, we focus on static modelling; in particular, we assume that the developer has already converted the relevant requirements into a preliminary class diagram such as that shown in Figure 1.
Figure 1: Library class model
The developer now wants to obtain a preliminary code that matches class diagram. Rather than using a UML diagram to code conversion tool, we will here rely on Copilot prompting with images. To achieve this, we attach Figure 1 and use the prompt illustrated in the following greyed box.
Implement the attached UML class diagram in python.
The resulting code is presented in our supplementary material (SM). Of course, depending on the specific prompt and GPT system used, the resulting code may vary. These variations are irrelevant, as SE typically proceeds via successive refinements to achieve the desired outcome.
To validate the classes generated against the original UML diagram, we can use Prompt Chaining and ask Copilot to generate a corresponding class diagram in PlantUML [^1] https://www.plantuml.com/ notation via the following prompt.
Develop a class diagram in PlantUML notation for the classes in the python code generated.
We can then use PlantUML to visualise the corresponding UML class diagram, which is shown in Figure 2 and, incidentally, only partly matches the original design in Figure 1. This is due to the fact that the Python implementation, whilst not incorrect, does not fully capture all cardinality constraints in the original diagram.
Figure 2: The reconstructed class diagram
4.2 Comparison to natural language
We consider once more the class diagram in Figure 1. To illustrate the advantage of using diagrams over natural language, consider the following prompt.
Describe the attached class diagram via natural language.
The result returned by Copilot is illustrated in Figure 3, this description is 1375 characters long.
The attached class diagram represents a library system with the following classes and relationships:
  1. 1.
    Classes:
    • Book: Represents a book in the library.
    • Copy: Represents a copy of a book or journal.
    • Journal: Represents a journal in the library.
    • LibraryMember: Represents a member of the library.
    • MemberOfStaff: Represents a staff member of the library.
  2. 2.
    Relationships:
    • Book to Copy: A book has a one-to-many relationship with copies, indicated by ”1” on the Book side and ”1..*” on the Copy side. This means one book can have multiple copies.
    • LibraryMember to Copy: A library member can borrow or return copies, indicated by ”0..1” on the LibraryMember side and ”0..*” on the Copy side. This means a library member can borrow or return zero or one copy, and a copy can be borrowed or returned by zero or more library members.
    • MemberOfStaff to LibraryMember: A member of staff is a type of library member, indicated by the inheritance arrow pointing from MemberOfStaff to LibraryMember.
    • MemberOfStaff to Journal: A member of staff can borrow or return journals, indicated by ”0..1” on the MemberOfStaff side and ”0..*” on the Journal side. This means a member of staff can borrow or return zero or one journal, and a journal can be borrowed or returned by zero or more staff members.
Figure 3: The textual description of the class diagram returned by Copilot
Contrast this with the PlantUML notation description of the class diagram illustrated in Figure 4, which is 302 characters long, that is only 20% of the size of the natural language description.
@startuml
class Book {
class Copy {
class Journal {
class LibraryMember {
class MemberOfStaff {
LibraryMember <|-- MemberOfStaff
LibraryMember "0..1" -- "0..*" Copy : borrows/returns
LibraryMember "0..1" -- "0..*" Journal : borrows/returns
Copy "1" -- "1..*" Book : is a copy of
@enduml
Figure 4: The description of the class diagram in PlantUML notation as returned by Copilot
This example makes apparent the point discussed in the paragraph from Tufte (2003) we quoted in Section 1: the visual channel is an intensely high-resolution channel.
4.3 Dynamic modelling
We may now want to proceed and describe the dynamic behaviour of the system. Rather than using natural language, we can describe specific aspect of such behaviour by leveraging suitable UML interaction diagrams, such as the sequence diagram in Figure 5.
Figure 5: A sequence diagram
To achieve this, we will attach Figure 5 and chain the following prompt in Copilot.
Implement the dynamic behaviour illustrated in the attached sequence diagram. Do not explicitly implement the actor “BookBorrower.”
The resulting code is presented in our SM.
Next, we may want to explore the behaviour of the various entities involved in a method call by obtaining a communication diagram for it. Copilot can generate this diagram (Figure 6) via the following prompt chained to previous outputs.
Generate a UML communication diagram in PlantUML notation illustrating the behaviour of the following code: member.borrow_copy(copy1).
Figure 6: A communication diagram
In the code generated by Copilot, the state of the book object changes when a copy of the book is successfully borrowed. In particular, the book may change from being borrowable (there is a copy of it in the library) to not borrowable (all copies are out on loan or reserved). This behaviour can be represented via a state diagram. Rather than drawing such state diagram, we generate one via the following prompt.
Generate a UML state diagram in PlantUML notation to represent the possible states of the Book object.
This leads to the state diagram in Figure 7.
Figure 7: The reconstructed state diagram
Assume now that the desired behaviour, illustrated in Figure 8, is slightly different from that obtained in this preliminary draft of our code.
Figure 8: The desired state diagram
We can ask Copilot to amend the code by chaining the following prompt.
Amend the python classes to capture the behaviour in the attached state diagram. Make sure the generated code continues to reflect the original class diagram.
The resulting code is presented in our SM.
Finally, we can ask Copilot to generate a new state diagram in PlantUML notation that reflects the behaviour of the updated code. The resulting diagram is shown in Figure 9 and matches the behaviour in Figure 8.
Figure 9: The revised state diagram
Unfortunately, the code obtained is not fully functional. There is a bug in method borrow_copy of class LibraryMember. This specific bug can be easily resolved by adding a single instruction (more details are provided in the SM). In general, further iterations of the process, potentially involving additional diagrams and a test-driven development strategy, can be used to obtain working code. In fact, the specific code generated or the absence/presence of bugs are both immaterial to the case study here presented, which is simply meant to illustrate potential use cases that combine multimodal GPTs and UML diagrams to innovate and support SE practices.
4.4 Hand drawn diagrams
While all the previous examples featured computer generated diagrams, Copilot is able to handle equally well hand-drawn diagrams. We consider the activity diagram shown in Figure 10.
Figure 10: An activity diagram
We attach the diagram, and input the following prompt.
Create a method in python that implements the attached UML activity diagram.
Copilot correctly recognises the fact that that the activity diagram represents a sorting algorithm, and returns a method implementing it.
Alternatively, assuming the code for our sorting algorithm is already available to us, we can ask Copilot to convert it to an activity diagram via the following prompt.
Create an activity diagram in PlantUML notation for the following method in python: [insert python code].
In this specific instance, Copilot returned a diagram with a small number of syntax errors. The errors were minor issues with the PlantUML syntax of the two while loops in the diagram, and could be easily fixed by hand. The resulting activity diagram is shown in Figure 11. Conversely, a second prompt asking Copilot to generate a diagram in PlantUML notation by translating directly the diagram in Figure 10 produced no errors. Once more, the focus of our discussion is facilitating, and not fully automating, SE activities. Therefore, the presence of minor errors is of no concern.
Figure 11: A PlantUML activity diagram of a sorting algorithm
4.5 Design patterns
Design Patterns Gamma et al. (1994) are reusable solutions to commonly occurring problems in software design. They are not specific implementations but rather general templates that can be adapted to various situations. By using design patterns, developers can create more flexible, maintainable, and efficient software systems. There exists a wealth of existing patterns available in the SE literature; In addition, new patterns can be created and illustrate via appropriate UML diagrams. For instance, we may consider the “Adapter” pattern, which converts the interface of a class into another interface clients expect. This pattern is illustrated in Figure 12.
Figure 12: The “Adapter” Pattern
Design patterns offer endless opportunities to carry out Few-Shot and CoT prompting. We shall illustrate this point with a practical example involving the “Adapter” pattern.
Consider the class diagram in Figure 13.
Figure 13: The data transfer system
In this system we have two portable drives, a Samsung drive and a Western Digital drive. The Samsung drive features a microusb connector, while the Western Digital drive features a traditional usb connector. Our aim is to allow the Samsung drive to read and write data via usb. To achieve this, we attach the diagram to Copilot and provide the following prompt.
Extend the UML diagram by using the Adapter design patter to allow a Samsung drive to read and write data via Usb. Provide the output in PlantUML notation.
The resulting diagram, which correctly implements the Adapter pattern, is illustrated in Figure 14.
Figure 14: The data transfer system with an adaptor.
Since the Adapter pattern is well known, and there is no need to provide a diagram for it while prompting. However, if the user intends to utilise a brand new pattern, then Few-Shot prompting can be leveraged to obtain the desired result. In practice, this would entail attaching a UML diagram representing the new design pattern that needs to be operationalised in the context of the given coding task.
5 Discussion
Throughout Section 4 we have utilised Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments. To ensure reproducibility, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in our SM.
In this section, we reflect on how the use cases discussed in Section 4, which are made possible thanks to the enhanced capabilities of GPT-4, address several research gaps in GenSE, as well as some long standing issues in SE.
We shall focus first on the research gaps we bridged in GenSE. First, our study addresses the lack of applications of GenAI to the software design sub-task of SE, a literature gap identified in Huang et al. (2024); Hou et al. (2024). Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts. Third, we have discussed and illustrated with an empirical example the superiority, in terms of information transfer efficiency, of diagram-based prompting against pure natural language prompting; a result that aligns with the discussion in Tufte (2003). Fourth, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries. Fifth, while there exist studies that investigated the production of PlantUML code from hand drawn diagrams Conrardy and Cabot (2024); Wang et al. (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation. Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature.
Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work. In Baltes and Diehl (2014), the authors speculated that, as documentation is frequently poorly written and out of date, sketches may serve as a supplement to conventional documentation like source code comments. Consequently, they advocated the development of suitable tools to support archiving and retrieval of such sketches. No such tool has been developed ever since and poorly documented code remains a persistent issue in SE. However, GenAI has already started to change this picture via automatic generation of code comments. And yet, if sketches and diagram (both formal and informal) can be transformed, as we have shown, into a preliminary draft of the desired source code, then producing reliable visual descriptions of the software system under development suddenly becomes appealing. Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding and boost team communication.

### v2

@@
 Similar conclusions are reached in Huang et al.
-(2024), which focused on 7 sub-tasks of SE: requirements generation, code generation, code summarisation, test generation, patch generation, code optimisation, and code translation.
-In all cases, the pipeline typically begins with instructions in natural language that need to be used in the context of one or more of these 7 sub-tasks.
+(2024), which focused on seven sub-tasks of SE: requirements generation, code generation, code summarisation, test generation, patch generation, code optimisation, and code translation.
+In all cases, the pipeline typically begins with instructions in natural language that need to be used in the context of one or more of these seven sub-tasks.
 In SE it is often the case that communication among developers, and between developers and customers, occurs in the form of sketches and diagrams Baltes and Diehl (2014) and not just via natural language.
@@
 The rest of this paper is organised as follows: Section 2 provides relevant background in GenAI and SE.
-Section 3 surveys contemporary academic literature on role of GenAI in SE.
+Section 3 outlines the methodology adopted in our study.
 Section 4 investigates the role of multimodal GPTs in software development by illustrating a selection of use cases.
-Section 5 provides some concluding remarks by reflecting on how the former use cases address several research gaps in Generative Software Engineering (GenSE), as well as some long standing issues in SE.
+Section 5 provides concluding remarks by reflecting on how the former use cases address several research gaps in Generative Software Engineering (GenSE), as well as some longstanding issues in SE, and open new research directions.
 2 Background
 This section provides relevant background in GenAI and SE.
-2.1 Software Engineering
-The term “software engineering” was coined by Margaret Hamilton, who developed on-board flight software for NASA’s Apollo program.
-Key milestones in the field include the development of the Waterfall model (1970s), the introduction of Object-Oriented programming (1980s), the Agile Manifesto (2001), which advocates for flexible, iterative development, the rise of DevOps (2000s) and of continuous integration/continuous delivery (CI/CD) pipelines (2010s).
-In the original Waterfall model, the software development process proceeds through a series of sequential phases: requirement analysis; software design; software development; software quality assurance; and software maintenance.
-To address the inflexibility of the Waterfall model, the increasing complexity of software, and poor customer satisfaction, in the early 2000s a set of frameworks and practices were proposed based on the values and principles expressed in the Manifesto for Agile Software Development Beck et al.
-(2001).
-Core values and principles of Agile include: favouring individuals and interactions over processes and tools; working software over comprehensive documentation; customer collaboration over contract negotiation; and responding to change over following a plan.
-Agile methodologies promote iterative development, frequent feedback, and adaptability to changing requirements throughout the aforementioned phases.
-At the same time, the Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al.
+2.1 Software Engineering and UML
+The Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al.
 (2004) utilised across the phases of the software development process.
@@
 In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system.
-Use case development (or development of user stories in Agile) is a key step of requirement analysis.
-Use cases describe how a user interacts with a system or product to achieve a specific goal.
+Use case development is a key step of requirement analysis.
+Use cases [^1] User stories in Agile Software Development Beck et al.
+(2001) describe how a user interacts with a system or product to achieve a specific goal.
 Static analysis of a software system describes the structure of a system by showing its classes, their attributes, operations (or methods), and the relationships among objects.
@@
 However, it should be noted that UML is a vast language that finds countless applications to go beyond the three areas here considered.
-Despite UML being the standard visual language for designing systems, recent studies suggest that that most software developers favour informal hand drawn diagrams and do not use UML; those using UML, tend to use it informally and selectively Baltes and Diehl (2014).
+Despite UML being the standard visual language for designing systems, recent studies suggest that that most software developers favour informal hand-drawn diagrams and do not use UML; those using UML, tend to use it informally and selectively Baltes and Diehl (2014).
 The authors further advocated the development of suitable tools to make better use of such sketches.
@@
 (2020).
-The latest GPTs, such as GPT-4 OpenAI (2024), are multimodal: they accept image and text inputs and produce text outputs, leading to new applications Li et al.
-(2024b).
-2.2.1 Prompt engineering
-Prompts are functions that map an example from a dataset to a natural language input and target output Bach et al.
-(2022).
-Due to the cost of re-training LLMs, prompt engineering, that is the use of prompts to train and query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning.
-Key prompt engineering techniques include: Single Prompt Techniques, Multiple Prompt Techniques, and the use of LLM frameworks with external tools.
-Single Prompt Techniques include Zero-Shot, Few-Shot, Chain of Thought (CoT), and Program-Aided Language Models (PAL).
-Zero-Shot Prompting involves providing tasks with natural language without additional context, relying on the model’s pre-existing capabilities.
-Few-Shot Prompting includes providing examples within the prompt to guide the model, enhancing its performance on complex tasks by exposing it to the input and output patterns Brown et al.
-(2020).
-Chain of Thought Prompting Wei et al.
-(2024) aids in breaking down reasoning tasks into smaller steps to improve processing and outcome accuracy, using either zero-shot or few-shot methods to encourage step-by-step thinking.
-PAL extends this by integrating code into reasoning processes to produce more accurate outputs Gao et al.
-(2023).
-Multiple Prompt Techniques leverage combining different strategies to refine outputs.
-Voting/self-consistency Wang et al.
-(2023) involves generating multiple responses and selecting the most common result, which can improve accuracy, especially for complex reasoning tasks.
-Divide and Conquer methods split tasks into subtasks handled in sequence for improved manageability and precision, seen in Directional Stimulus Prompting Li et al.
-(2024a), Generated Knowledge Liu et al.
-(2021), and Prompt Chaining.
-Self-evaluation asks the model to verify output accuracy, exemplified by Reflexion Shinn et al.
-(2024) and Tree of Thoughts Yao et al.
-(2024), enabling iterative improvement.
-Retrieval-Augmented Generation (RAG) Fan et al.
-(2024) and ReAct Yao et al.
-(2023) combine LLMs with external systems to improve context handling and output relevance.
-RAG uses vector search for context retrieval before passing it to an LLM for output generation.
-ReAct combines reasoning traces and task-specific actions, allowing models to interact with external knowledge sources to refine outputs.
-3 The role of GenAI in Software Engineering
-In recent times, SE has become one of the important application areas for GenAI. We focus on two recent surveys Huang et al. (2024); Hou et al. (2024) investigating the interplay between SE and GenAI.
+The latest GPTs, such as GPT-4 OpenAI (2024), are multimodal: they accept image and text inputs, and produce image and text outputs, leading to new applications.
+Prompts are instructions or input given to a large language model (LLM) to generate a specific response.
+Prompt engineering, that is the design prompting strategies to query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning.
+Prompt engineering techniques include: Single Prompt Techniques (e.g.
+Zero-Shot, Few-Shot), Multiple Prompt Techniques (e.g.
+Prompt Chaining), and the use of LLM frameworks with external tools (e.g.
+Retrieval-Augmented Generation Fan et al.
+(2024), and ReAct Yao et al.
+(2023)).
+2.3 The role of GenAI in Software Engineering
+In recent times, SE has become one of the important application areas for GenAI.
+We focus on two recent surveys Huang et al.
+(2024); Hou et al.
+(2024) investigating the interplay between SE and GenAI over a large body of recent works.
 Several studies, e.g.
@@
 (2024), investigated the use of GenAI in the context of the requirement engineering sub-task.
-We do not believe sketches and diagram play a significant role in this sub-task, which is mainly concerned in turning requirements expressed in natural language by the customer into suitable user stories and/or conceptual diagrams Robeer et al.
-(2016).
-SE sub-tasks of interest in the context of the present study include: software design, software development, and code summarisation.
+GenAI does play a role in this sub-task, which is mainly concerned in turning requirements expressed in natural language by the customer into suitable user stories and/or conceptual diagrams Robeer et al.
+(2016), but our focus in this study will not be on this sub-task.
+Conversely, SE sub-tasks of interest in the context of the present study include: software design, software development, and code summarisation.
 Application of LLMs in software design remains relatively sparse: Huang et al.
-(2024) does not cover this sub-task, while Hou et al.
+(2024) does not include any study within this sub-task; while Hou et al.
 (2024) only report 4 works Kolthoff et al.
@@
 (2024); Zhang et al.
-(2024), none of which overlaps with the content of the present study; they also stress that by expanding the use of LLMs to this under-explored area, it is possible to improve how software designs are conceptualised.
-Conversely, both surveys identify a plethora of works concerned with software development (including code generation, test case generation, patch generation, and code optimisation) and code summarisation.
+(2024), none of which overlaps with the content of the present study; they also stress that by expanding the use of LLMs to this under-explored area it is possible to improve how software designs are conceptualised.
+Both surveys identify a plethora of works concerned with software development (including code generation, test case generation, patch generation, and code optimisation) and code summarisation.
 Code generation has been object of investigation for a long time in the AI community.
@@
 This capability has significantly advanced code completion, automatic code generation, and the transformation of natural language annotations into code, empowering developers with powerful tools and driving automation in software development.
-None of the studies listed in the above surveys focus on code generation that leverages multimodal prompts that include UML diagrams.
+None of the studies listed in the above surveys focus on code generation leveraging multimodal prompts that include sketches & diagrams.
 Code summarisation Ahmed et al.
@@
 Existing studies in code summarisation focus on analysing code structures and contexts to generate informative natural language summaries.
-None of the studies listed in the above surveys focus on code summarisation producing a UML diagram as its output.
+None of the studies listed in the above surveys focus on code summarisation producing a diagram as its output.
 Finally, while there exist a few studies that investigated the generation of UML diagram with support from LLMs Conrardy and Cabot (2024); Wang et al.
@@
 Perhaps the most interesting study among the three listed is Cámara et al.
-(2023), which focus on how to build UML class diagrams by using ChatGPT as a modelling assistant by leveraging the PlantUML language, which we will also adopt in the following sections.
+(2023), which focuses on building UML class diagrams in PlantUML notation, which we also adopt in this work, by using ChatGPT as a modelling assistant prompted with instructions in natural language.
+3 Methodology
+Our approach takes the form of an exploratory case study Priya (2020), which aims to identify new research questions that can be investigated in future studies.
+More specifically, we develop a portfolio of novel GenSE use cases that, to the best of our knowledge, have not been previously investigated in the literature.
+In our study, we utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
+To ensure reproducibility of the discussion in Sections 4.1-4.5, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in the supplementary material (SM).
+In both cases, we left all LLM parameters to their default settings and we did not specify any role or context instructions.
+Table 1 maps use cases discussed in the rest of this work to relevant SE sub-tasks of interest.
+All use cases are based on an interaction with the LLMs that takes the form of an interactive chat comprising multiple rounds of questions and responses, which allow users to step incrementally toward answers and thus get help with multipart problems.
+The core principle underpinning all our use cases is to leverage UML diagrams in order to guide the software development process.
+More specifically, in Section 4.1 we leverage class diagrams to guide the LLM in the context of implementing relevant classes, attributes, and operations; in Section 4.2, we contrast diagram-based and natural language-based descriptions of a system; in Section 4.3 we leverage interaction diagrams to guide the LLM in the context of implementing the desired system behaviour; in Section 4.4 we leverage hand drawn activity diagrams to guide the LLM in the context of implementing a given method; in Section 4.5 we leverage design patterns to influence the design of a given system.
+Finally, in our SM we present three additional case studies: in the first case study, we leverage hand drawn diagrams and design patterns to implement a mathematical expressions evaluator; in the second and third case studies, which feature a higher degree of complexity, we leverage class diagrams to implement a tic tac toe game and a game of checkers, respectively.
+Use case Section SE sub-tasks
+static modelling 4.1 SD/DE/CS
+information transfer 4.2 SD
+dynamic modelling 4.3 SD/DE/CS
+hand drawn diagrams 4.4 DE/CS
+design patterns 4.5 SD
+expression evaluator SM SD/DE
+tic tac toe SM SD/DE
+checkers SM SD/DE
+Table 1: Use cases investigated in the rest of this work (SD: software design; DE: software development; CS: code summarisation)
 4 Multimodal GPTs in software development
 In what follows, we will assume that a preliminary requirement analysis has been carried out, which has produced a portfolio of initial user stories.
-While GPTs may find applications in the realm of the development of user stories and in the translation of user stories into working software, as things stand today, this level of automation in the realm of code generation does not exist in SE practice.
-Conversely, what we find in practice are teams of software developers tasked with translating user stories into working software.
-Our focus in this section, is to explore novel use cases of GPTs in this specific context.
+While GPTs may in principle support automated elicitation of user stories and fully automated translation of user stories into working software, as things stand today this level of automation in the realm of code generation is hardly found in SE practice; nor we suspect it would be beneficial, as it may conflict with some of the principles of Agile.
+What we find in practice are teams of software developers collaborating to translate user stories into working software.
+Our focus is to illustrate novel use cases of GPTs in this specific context.
 The role of sketches and diagrams in the daily work of software developers has been investigated in Baltes and Diehl (2014).
-This study found that most practitioners produced informal hand drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively.
+This study found that most practitioners produced informal hand-drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively.
 We argue that this stems from the fact that the adoption of a “formal” notation, at present, bears no advantage with respect to an “informal” one.
@@
 While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools.
-We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one for generating complete code.
+We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one aimed at describing a complete system.
 In other words, UML diagrams are most useful when they are high level and sufficiently abstract.
 Generation of complete working code requires such diagrams to reach a level of detail that is equivalent to the working code itself; but generating diagrams at this level of abstraction would be a complete waste of time: why then not generating the code itself directly?
-Albeit we believe there is a good reason for practitioners to produce informal hand drawn diagrams, we also believe that this practice is set to change with the advent of multimodal GPTs, such as GPT-4 OpenAI (2024).
-In what follows, we will outline a set of use cases illustrating how practitioners may combine multimodal GPTs and UML diagrams to innovate SE practices.
-We will utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
+Albeit there are already good reasons for practitioners to produce sketches and diagrams in certain circumstances, we believe the advent of multimodal GPTs, such as GPT-4 OpenAI (2024), will substantially increase the associated use cases.
+To exemplify this, in what follows we outline a set of use cases illustrating how practitioners may combine multimodal GPTs and UML diagrams to innovate SE practices.
 4.1 Static modelling
-As a motivating example to illustrate our use cases, we consider the introductory case study in Section 3 of Stevens and Pooley (2006).
+As a motivating example to illustrate our use cases, we consider the introductory case study in Chapter 3 of Stevens and Pooley (2006).
 In this section, we focus on static modelling; in particular, we assume that the developer has already converted the relevant requirements into a preliminary class diagram such as that shown in Figure 1.
-Figure 1: Library class model
-The developer now wants to obtain a preliminary code that matches class diagram.
+Figure 1: Class model of a library
+The developer now wants to obtain a preliminary code that matches this class diagram.
 Rather than using a UML diagram to code conversion tool, we will here rely on Copilot prompting with images.
@@
 Implement the attached UML class diagram in Python.
-The resulting code is presented in our supplementary material (SM).
-Of course, depending on the specific prompt and GPT system used, the resulting code may vary.
-These variations are irrelevant, as SE typically proceeds via successive refinements to achieve the desired outcome.
+The resulting code is presented in the SM.
+It is noteworthy that in addition to implementing the classes, Copilot also made an attempt at drafting part of the business logic by leveraging semantic information associated with the labels of the relationships in the diagram.
 To validate the classes generated against the original UML diagram, we can use Prompt Chaining and ask Copilot to generate a corresponding class diagram in PlantUML [^2] https://www.plantuml.com/ notation via the following prompt.
@@
 @enduml
-Figure 4: The description of the class diagram in PlantUML notation as returned by Copilot
+Figure 4: The description of the class diagram in PlantUML notation
 This example makes apparent the point discussed in the paragraph from Tufte (2003) we quoted in Section 1: the visual channel is an intensely high-resolution channel.
+Moreover, it reinforces the findings in Li et al.
+(2024) on the interrelation between understanding, learning, and compression.
+Further examples contrasting diagrams and natural language descriptions are provided in the SM.
 4.3 Dynamic modelling
@@
 Rather than using natural language, we can describe specific aspect of such behaviour by leveraging suitable UML interaction diagrams, such as the sequence diagram in Figure 5.
-Figure 5: A sequence diagram
+Figure 5: A sequence diagram illustrating an interaction
 To achieve this, we will attach Figure 5 and chain the following prompt in Copilot.
 Implement the dynamic behaviour illustrated in the attached sequence diagram. Do not explicitly implement the actor “BookBorrower.”
-The resulting code is presented in our SM.
+The resulting code is presented in the SM.
 Next, we may want to explore the behaviour of the various entities involved in a method call by obtaining a communication diagram for it.
@@
 Amend the Python classes to capture the behaviour in the attached state diagram. Make sure the generated code continues to reflect the original class diagram.
-The resulting code is presented in our SM.
+The resulting code is presented in the SM.
 Finally, we can ask Copilot to generate a new state diagram in PlantUML notation that reflects the behaviour of the updated code.
@@
 There is a bug in method borrow_copy of class LibraryMember.
-This specific bug can be easily resolved by adding a single instruction (more details are provided in the SM).
-In general, further iterations of the process, potentially involving additional diagrams and a test-driven development strategy, can be used to obtain working code.
-In fact, the specific code generated or the absence/presence of bugs are both immaterial to the case study here presented, which is simply meant to illustrate potential use cases that combine multimodal GPTs and UML diagrams to innovate and support SE practices.
-4.4 Hand drawn diagrams
+This bug can easily be resolved by adding a single instruction (more details are provided in the SM).
+In line with contemporary SE and Agile practices — which predicate an iterative form of software development aimed at progressively discovering how to meet customer requirements — further iterations of the process, potentially involving additional diagrams and further prompts, can be used to obtain working code.
+4.4 Hand-drawn diagrams
 While all the previous examples featured computer generated diagrams, Copilot is able to handle equally well hand-drawn diagrams. We consider the activity diagram shown in Figure 10.
@@
 Create a method in Python that implements the attached UML activity diagram.
-Copilot correctly recognises the fact that that the activity diagram represents a sorting algorithm, and returns a method implementing it.
+Copilot correctly recognises the fact that that the activity diagram represents a sorting algorithm (bubble sort), and returns a method implementing it.
 Alternatively, assuming the code for our sorting algorithm is already available to us, we can ask Copilot to convert it to an activity diagram via the following prompt.
@@
 By using design patterns, developers can create more flexible, maintainable, and efficient software systems.
-There exists a wealth of existing patterns available in the SE literature; In addition, new patterns can be created and illustrate via appropriate UML diagrams.
+There exists a wealth of existing patterns available in the SE literature; In addition, new patterns can be created and illustrated via appropriate UML diagrams.
 For instance, we may consider the “Adapter” pattern, which converts the interface of a class into another interface clients expect.
@@
 Design patterns offer endless opportunities to carry out Few-Shot and CoT prompting. We shall illustrate this point with a practical example involving the “Adapter” pattern.
+Figure 13: The data transfer system
 Consider the class diagram in Figure 13.
-Figure 13: The data transfer system
 In this system we have two portable drives, a Samsung drive and a Western Digital drive.
@@
 In practice, this would entail attaching a UML diagram representing the new design pattern that needs to be operationalised in the context of the given coding task.
-5 Discussion
-Throughout Section 4 we have utilised Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
-To ensure reproducibility, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in our SM.
+5 Discussion and conclusions
 In this section, we reflect on how the use cases discussed in Section 4, which are made possible thanks to the enhanced capabilities of GPT-4, address several research gaps in GenSE, as well as some long standing issues in SE.
+We will also discuss how the same use cases open new research directions that can be investigated in future studies.
 We shall focus first on the research gaps we bridged in GenSE.
@@
 Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts.
-Third, we have discussed and illustrated with an empirical example the superiority, in terms of information transfer efficiency, of diagram-based prompting against pure natural language prompting; a result that aligns with the discussion in Tufte (2003).
+Third, we have discussed and illustrated with an empirical example that diagram-based prompting may be advantageous, in terms of information transfer efficiency, compared to pure natural language prompting; a result that aligns with the discussion in Tufte (2003).
 Fourth, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries.
-Fifth, while there exist studies that investigated the production of PlantUML code from hand drawn diagrams Conrardy and Cabot (2024); Wang et al.
+Fifth, while there exist studies that investigated the production of PlantUML code from hand-drawn diagrams Conrardy and Cabot (2024); Wang et al.
 (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation.
-Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature.
+Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature, and hence merits attention in future studies.
 Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work.
@@
 And yet, if sketches and diagram (both formal and informal) can be transformed, as we have shown, into a preliminary draft of the desired source code, then producing reliable visual descriptions of the software system under development suddenly becomes appealing.
-Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding and boost team communication.
+Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding, boost team communication, and resolve the long lasting conflict between investing time in developing working software or writing comprehensive documentation.
+These are all important and novel directions of enquiry that should be investigated in future studies.

### v3

@@
 2.1 Software Engineering and UML
-The Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al.
-(2004) utilised across the phases of the software development process.
-UML is a general-purpose standardised visual language for designing systems.
+The Unified Modelling Language (UML) is a general-purpose standardised visual language for designing systems.
 In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system.
@@
 Prompts are instructions or input given to a large language model (LLM) to generate a specific response.
-Prompt engineering, that is the design prompting strategies to query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning.
-Prompt engineering techniques include: Single Prompt Techniques (e.g.
-Zero-Shot, Few-Shot), Multiple Prompt Techniques (e.g.
-Prompt Chaining), and the use of LLM frameworks with external tools (e.g.
-Retrieval-Augmented Generation Fan et al.
-(2024), and ReAct Yao et al.
-(2023)).
+Prompt engineering — the design prompting strategies to query language models — offers a cost-effective way to adapt pre-trained models without full fine-tuning.
+Single Prompt Techniques.
+Zero-Shot Prompting involves providing tasks with natural language without additional context, relying on the model’s pre-existing capabilities.
+Few-Shot Prompting includes providing examples within the prompt to guide the model, enhancing its performance on complex tasks by exposing it to the input and output patterns Brown et al.
+(2020).
+Chain of Thought Prompting Wei et al.
+(2024) aids in breaking down reasoning tasks into smaller steps to improve outcome accuracy, using either zero-shot or few-shot methods to encourage step-by-step thinking.
+Multiple Prompt Techniques.
+Voting/self-consistency Wang et al.
+(2023) involves generating multiple responses and selecting the most common result, which can improve accuracy, especially for complex reasoning tasks.
+Divide and Conquer methods split tasks into subtasks handled in sequence for improved manageability and precision, seen in Directional Stimulus Prompting Li et al.
+(2024a), Generated Knowledge Liu et al.
+(2021), and Prompt Chaining.
+Self-evaluation asks the model to verify output accuracy, exemplified by Reflexion Shinn et al.
+(2024) and Tree of Thoughts Yao et al.
+(2024), enabling iterative improvement.
+Retrieval-Augmented Generation (RAG) Fan et al. (2024) and ReAct Yao et al. (2023) combine LLMs with external systems to improve context handling and output relevance.
 2.3 The role of GenAI in Software Engineering
@@
 (2023), can generate code directly from natural language descriptions.
-This capability has significantly advanced code completion, automatic code generation, and the transformation of natural language annotations into code, empowering developers with powerful tools and driving automation in software development.
-None of the studies listed in the above surveys focus on code generation leveraging multimodal prompts that include sketches & diagrams.
+While there are several works and benchmarks in the literature concerned with method-level code generation, to the best of our knowledge there is only one study and benchmark on class-level code generation Du et al.
+(2023), and none on diagram-level code generation.
+Moreover, none of the studies listed in the above surveys focus on code generation leveraging multimodal prompts that include sketches & diagrams.
 Code summarisation Ahmed et al.
@@
 3 Methodology
-Our approach takes the form of an exploratory case study Priya (2020), which aims to identify new research questions that can be investigated in future studies.
-More specifically, we develop a portfolio of novel GenSE use cases that, to the best of our knowledge, have not been previously investigated in the literature.
-In our study, we utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
-To ensure reproducibility of the discussion in Sections 4.1-4.5, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in the supplementary material (SM).
+We develop a portfolio of novel GenSE use cases that, to the best of our knowledge, have not been previously investigated in the literature.
+We utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
+To ensure reproducibility of the discussion in Sections 4.1-4.4, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in the supplementary material (SM).
 In both cases, we left all LLM parameters to their default settings and we did not specify any role or context instructions.
 Table 1 maps use cases discussed in the rest of this work to relevant SE sub-tasks of interest.
-All use cases are based on an interaction with the LLMs that takes the form of an interactive chat comprising multiple rounds of questions and responses, which allow users to step incrementally toward answers and thus get help with multipart problems.
-The core principle underpinning all our use cases is to leverage UML diagrams in order to guide the software development process.
-More specifically, in Section 4.1 we leverage class diagrams to guide the LLM in the context of implementing relevant classes, attributes, and operations; in Section 4.2, we contrast diagram-based and natural language-based descriptions of a system; in Section 4.3 we leverage interaction diagrams to guide the LLM in the context of implementing the desired system behaviour; in Section 4.4 we leverage hand drawn activity diagrams to guide the LLM in the context of implementing a given method; in Section 4.5 we leverage design patterns to influence the design of a given system.
-Finally, in our SM we present three additional case studies: in the first case study, we leverage hand drawn diagrams and design patterns to implement a mathematical expressions evaluator; in the second and third case studies, which feature a higher degree of complexity, we leverage class diagrams to implement a tic tac toe game and a game of checkers, respectively.
+All use cases are based on a duly documented interaction with the LLMs that takes the form of a chat comprising multiple rounds of questions and responses (Prompt Chaining), which allow users to step incrementally towards answers and thus get help with multipart problems Google AI (2025).
+The core principle underpinning all our use cases is to illustrate possible strategies to leverage UML diagrams in order to guide the software development process.
+More specifically, in Section 4.1 we leverage class diagrams to guide the LLM in the context of implementing relevant classes, attributes, and operations; in Section 4.2 we leverage interaction diagrams to guide the LLM in the context of implementing the desired system behaviour; in Section 4.3 we leverage hand-drawn activity diagrams to guide the LLM in the context of implementing a given method; in Section 4.4 we leverage design patterns to influence the design of a given system.
+Finally, in Section 4.5 we present three additional case studies: in the first, we leverage hand-drawn diagrams and design patterns to implement a mathematical expressions evaluator; in the second and third, which feature a higher degree of complexity, we leverage class diagrams to implement a tic tac toe game and a game of checkers, respectively.
 Use case Section SE sub-tasks
 static modelling 4.1 SD/DE/CS
-information transfer 4.2 SD
-dynamic modelling 4.3 SD/DE/CS
-hand drawn diagrams 4.4 DE/CS
-design patterns 4.5 SD
-expression evaluator SM SD/DE
-tic tac toe SM SD/DE
-checkers SM SD/DE
+dynamic modelling 4.2 SD/DE/CS
+hand drawn diagrams 4.3 DE/CS
+design patterns 4.4 SD
+expression evaluator 4.5 SD/DE
+tic tac toe 4.5 SD/DE
+checkers 4.5 SD/DE
 Table 1: Use cases investigated in the rest of this work (SD: software design; DE: software development; CS: code summarisation)
@@
 Over several decades, firms have repeatedly tried to develop tools (e.g.
-IBM Rational Rose) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code.
+IBM Rational) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code.
 While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools.
-We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one aimed at describing a complete system.
+The lacklustre success of these tools is likely due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one aimed at describing a complete system.
 In other words, UML diagrams are most useful when they are high level and sufficiently abstract.
@@
 Figure 2: The reconstructed class diagram
-4.2 Comparison to natural language
-We consider once more the class diagram in Figure 1. To illustrate the advantage of using diagrams over natural language, consider the following prompt.
-Describe the attached class diagram via natural language.
-The result returned by Copilot is illustrated in Figure 3, this description is 1375 characters long.
-The attached class diagram represents a library system with the following classes and relationships:
-  1. 1.
-    Classes:
-    • Book: Represents a book in the library.
-    • Copy: Represents a copy of a book or journal.
-    • Journal: Represents a journal in the library.
-    • LibraryMember: Represents a member of the library.
-    • MemberOfStaff: Represents a staff member of the library.
-  2. 2.
-    Relationships:
-    • Book to Copy: A book has a one-to-many relationship with copies, indicated by ”1” on the Book side and ”1..*” on the Copy side. This means one book can have multiple copies.
-    • LibraryMember to Copy: A library member can borrow or return copies, indicated by ”0..1” on the LibraryMember side and ”0..*” on the Copy side.
-This means a library member can borrow or return zero or one copy, and a copy can be borrowed or returned by zero or more library members.
-    • MemberOfStaff to LibraryMember: A member of staff is a type of library member, indicated by the inheritance arrow pointing from MemberOfStaff to LibraryMember.
-    • MemberOfStaff to Journal: A member of staff can borrow or return journals, indicated by ”0..1” on the MemberOfStaff side and ”0..*” on the Journal side.
-This means a member of staff can borrow or return zero or one journal, and a journal can be borrowed or returned by zero or more staff members.
-Figure 3: The textual description of the class diagram returned by Copilot
-Contrast this with the PlantUML notation description of the class diagram illustrated in Figure 4, which is 302 characters long, that is only 20% of the size of the natural language description.
-@startuml
-class Book {
-class Copy {
-class Journal {
-class LibraryMember {
-class MemberOfStaff {
-LibraryMember <|-- MemberOfStaff
-LibraryMember "0..1" -- "0..*" Copy : borrows/returns
-LibraryMember "0..1" -- "0..*" Journal : borrows/returns
-Copy "1" -- "1..*" Book : is a copy of
-@enduml
-Figure 4: The description of the class diagram in PlantUML notation
-This example makes apparent the point discussed in the paragraph from Tufte (2003) we quoted in Section 1: the visual channel is an intensely high-resolution channel.
-Moreover, it reinforces the findings in Li et al.
-(2024) on the interrelation between understanding, learning, and compression.
-Further examples contrasting diagrams and natural language descriptions are provided in the SM.
 4.2 Dynamic modelling
 We may now want to proceed and describe the dynamic behaviour of the system.
-Rather than using natural language, we can describe specific aspect of such behaviour by leveraging suitable UML interaction diagrams, such as the sequence diagram in Figure 5.
+Rather than using natural language, we may describe specific aspect of such behaviour by leveraging suitable UML interaction diagrams, such as the sequence diagram in Figure 3.
 Figure 3: A sequence diagram illustrating an interaction
@@
 Figure 7: The revised state diagram
-Unfortunately, the code obtained is not fully functional.
-There is a bug in method borrow_copy of class LibraryMember.
-This bug can easily be resolved by adding a single instruction (more details are provided in the SM).
-In line with contemporary SE and Agile practices — which predicate an iterative form of software development aimed at progressively discovering how to meet customer requirements — further iterations of the process, potentially involving additional diagrams and further prompts, can be used to obtain working code.
 4.3 Hand-drawn diagrams
@@
 Conversely, a second prompt asking Copilot to generate a diagram in PlantUML notation by translating directly the diagram in Figure 8 produced no errors.
-Once more, the focus of our discussion is facilitating, and not fully automating, SE activities.
-Therefore, the presence of minor errors is of no concern.
+Since the focus of our discussion is facilitating — and not fully automating — SE activities, the presence of minor errors is of no concern, as these can easily be addressed in follow up prompts.
 Figure 9: A PlantUML activity diagram of a sorting algorithm
@@
 In practice, this would entail attaching a UML diagram representing the new design pattern that needs to be operationalised in the context of the given coding task.
+4.5 Implementation of complex systems
+In this section, we discuss the implementation of three more complex applications.
+In all cases, the interaction takes the form of a chat comprising multiple rounds of questions and responses (Prompt Chaining), which allows the user to step incrementally towards a working implementation.
+4.5.1 Evaluating mathematical expressions
+In this first application, we consider a class diagram (provided in our SM) that conceptually captures a system that evaluates mathematical expressions comprising additions and subtractions.
+The system comprises five interrelated classes.
+We were able to obtain a complete implementation with a single multimodal prompt (diagram + instruction to implement the diagram).
+An additional prompt was required to implement an evaluator by leveraging the Visitor design pattern.
+4.5.2 Tic tac toe
+We consider the class diagram [^3] https://github.com/pelensky/JavaTTT/blob/master/UML.pdf for a tic tac toe game.
+The diagram comprises over twenty interrelated classes.
+We use this diagram to guide the development of a tic tac toe Python application.
+The detailed interaction with Copilot is illustrated in our SM.
+This interaction involves seven prompts in which Copilot is asked to fully implement specific classes in the diagram.
+In line with traditional SE practices, we start from classes featuring low complexity and connectivity, and we then move towards other classes that depend on those already implemented.
+This is followed by two prompts to correct errors encountered while trying to execute the application.
+In each of these prompts the full stack trace of the error is fed to Copilot.
+Finally, an additional prompt is used to tailor board visualisation.
+A complete application that neatly matches the conceptual model presented in the class diagram is hence obtained in ten short prompts comprising a single sentence each.
+4.5.3 Checkers
+We consider the class diagram [^4] https://app.genmymodel.com/api/repository/wanex505/checkers for a checkers game.
+The diagram comprises over ten interrelated classes and a complex underpinning logic.
+We use this diagram to guide the development of a checkers Python application.
+The detailed interaction with Copilot is illustrated in our SM.
+This interaction involves nine prompts in which Copilot is asked to fully implement specific classes in the diagram.
+This is followed by an additional prompt to develop a text-based drawing logic.
+Ten additional prompts are then required to correct errors encountered while trying to execute the application and play the game.
+These errors include both exceptions (e.g.
+a class attribute is missing) as well as conceptual errors that affect the game logic (a piece is not removed from the board after being jumped over) or the user interaction (e.g.
+the user is requested to input a move, but it is not stated if they are playing as black or white).
+A complete application that matches the conceptual model presented in the class diagram is hence obtained in twenty short prompts comprising a single sentence each.
+In our SM we provide a more detailed description of each case study, as well as Jupyter Notebooks reporting the complete interactions with Copilot.
 5 Discussion and conclusions
@@
 We will also discuss how the same use cases open new research directions that can be investigated in future studies.
-We shall focus first on the research gaps we bridged in GenSE.
+We shall next focus on the research gaps we bridged in GenSE.
 First, our study addresses the lack of applications of GenAI to the software design sub-task of SE, a literature gap identified in Huang et al.
@@
 Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts.
-Third, we have discussed and illustrated with an empirical example that diagram-based prompting may be advantageous, in terms of information transfer efficiency, compared to pure natural language prompting; a result that aligns with the discussion in Tufte (2003).
-Fourth, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries.
-Fifth, while there exist studies that investigated the production of PlantUML code from hand-drawn diagrams Conrardy and Cabot (2024); Wang et al.
+We argue that diagram-based prompting may be advantageous, in terms of information transfer efficiency, compared to pure natural language prompting, as diagrams leverage background knowledge associated with their visual elements and structure, thereby leading to compression Li et al.
+(2024b).
+Third, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries.
+Fourth, while there exist studies that investigated the production of PlantUML code from hand-drawn diagrams Conrardy and Cabot (2024); Wang et al.
 (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation.
-Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature, and hence merits attention in future studies.
+Moreover, to the best of our knowledge, this is the first study that focus on diagram-level (rather than method-level or class-level) code generation.
+Finally, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature, and hence merits attention in future studies.
 Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work.
-In Baltes and Diehl (2014), the authors speculated that, as documentation is frequently poorly written and out of date, sketches may serve as a supplement to conventional documentation like source code comments.
-Consequently, they advocated the development of suitable tools to support archiving and retrieval of such sketches.
-No such tool has been developed ever since and poorly documented code remains a persistent issue in SE.
+In Baltes and Diehl (2014), the authors suggested that sketches could supplement often outdated and poorly written documentation, advocating for tools to archive and retrieve these sketches.
+No such tool exists and poorly documented code remains a persistent issue in SE.
 However, GenAI has already started to change this picture via automatic generation of code comments.